# Learning Autonomy: Off-Road Navigation Enhanced by Human Input

### v1

@@ Abstract
I INTRODUCTION
@@
Our approach uses utility functions to directly extract key features from segmented images and learns human driving behaviour using demonstration data. This approach diverges from traditional methods, which typically require either extensive labeled data for end-to-end learning or precise sensor calibration and global mapping in classical robotics approaches. By focusing on extracting key features directly in the trajectory space, our method simplifies the process, avoiding the complexity of global map generation. This allows our algorithm to effectively learn driving patterns from as little as 5-10 minutes of driving data. Our approach reduces the need for large datasets and detailed calibration, facilitating a more straightforward and efficient learning algorithm for navigating complex environments. The main contributions of this work are outlined below:
@@
II Related Work
Classical work in off-road navigation [1, 2, 3, 4, 5] has focused on creating the costmap of the environment from sensors’ data to represent navigation cost associated with the various types of visual and geometric features of the environment. Earlier approaches [1, 3] relied on feature engineering while later approaches [6, 7] relied on deep learning based semantic segmentation to represent the visual and geometric features of the surrounding terrain. These later approaches train the semantic segmentation pipeline from scratch and use hand designed cost functions for the planner. Although, these costmaps provide rich information for the downstream planning tasks, tuning them to capture the complex dynamic interactions while navigating on various terrain types is extremely challenging and requires significant domain expertise.
Recent advances in Deep learning have inspired researchers in robotics community to develop end-to-end learning algorithms[8, 9] that directly learn the mapping from sensor information to control commands thereby bypassing the need for manual costmap creation and tuning. Despite the promises of these end-to-end approaches, they still require a large amount of data and show poor generalisation to different settings (domain adaptation). Moreover, the black box nature of these approaches make them very hard to debug and deploy on real systems.
@@ II Related Work
Recent works like [10, 11, 12, 13] leverage human driving data to either directly learn the costmap of the environment or learn the traversability of various terrain types. These costmaps are crucial to solve the optimal control problem and generate appropriate vehicle controls. While these methods effectively model the tire-terrain interactions of the vehicle, they often overlook the higher level reasoning required to navigate in challenging off-road environments. Additionally, these methods primarily depend on real-world data to learn costmaps and traversability. In contrast, our research demonstrates the potential of simulation platforms to learn a local planning algorithm.
III Problem Statement
@@ III-A Definitions
  • Preference set $S$: The preference set $S$ is a fixed set that represents all the different preferences (Trajectories) available for the human operator at any instant. To construct $S$ we first discretize the control commands—linear and angular velocities. Let $A=[\{v^{i},w^{i}\}_{i=1}^{i=m}]$, represent this fixed discrete set of linear and angular velocities. Assuming the vehicle starts at $[0,0,0]$, we then apply forward simulation using the differential drive kinematic model to generate $m$ distinct trajectories over a pre defined time horizon $n$.
@@
    $\theta^{i}_{t}$ $=w^{i}\theta^{i}_{t-1}\Delta t$
@@ IV METHODOLOGY
$U(S)$ $=\begin{bmatrix}u^{1}(T^{1})&u^{2}(T^{1})&\ldots&u^{5}(T^{1},S,P)\\ u^{1}(T^{2})&u^{2}(T^{3})&\ldots&u^{5}(T^{2},S,P)\\ \vdots&\vdots&\vdots&\vdots\\ u^{1}(T^{m})&u^{2}(T^{m})&\ldots&u^{5}(T^{m},S,P)\end{bmatrix}$ (1)
@@
Here $u^{1}(T),..,u^{4}(T)$ are utility functions associated with 4 different terrain features of the environment and only depend on a single trajectory. The utility function $u^{5}(T,S,P)$ represents the distance-based utility, which is dependent on a specific trajectory within the preference set $S$, the preference set as a whole, and the reference path $P$. In the next section we discuss these utility functions in detail.
IV-A Utility Functions
The utility feature defined in Eq. 1, consists of several utility functions each representing a specific feature of the environment. In this work, we consider five utility functions, which are categorized into terrain utility functions and distance utility functions. The first four, $u^{1}(T)$ through $u^{4}(T)$, are terrain utility functions that correspond to different terrain types in the environment: non-traversable, water, rock, and mud, respectively. The fifth utility function, $u^{5}(T,S,P)$, is a distance utility function, which measures a trajectory’s relative closeness to the reference path $P$ with respect to all the other trajectories in the preference set $S$. To calculate the utility functions corresponding to the 4 terrain types, we project the trajectory T on to the camera plane and then use the following equation.
@@ IV-A Utility Functions
The distance utility function for a trajectory is defined in the context of a preference set and a reference path. We assume that we are given a preference set $S=\{T^{i}\}_{i=1}^{i=m}$ and for each of these trajectories $T^{i}\in S$, the distance between the last point on the trajectory and the closest waypoint in the reference path to this last point is represented by $d(T^{i},P)$. Now we can define the distance utility function $(u^{5}(T^{i},S,P))$:
@@
Thus the distance utility function represents the relative closeness of a trajectory from the waypoints within a sample of trajectories.
IV-B Learning human preferences using demonstrations
To learn human preferences from human demonstrations we propose a supervised learning approach. Specifically, we frame the problem as a classification task where, at each time step, the planner selects a trajectory from a fixed preference set $S$ of trajectories. The labels for this classification task are derived from human demonstrations, where at each time step, the human operator selects a trajectory from the same fixed preference set $S$ while following a reference path. During this demonstration, the human operator follows the reference path as closely as possible, while simultaneously avoiding or preferring different terrain types based on the real-time camera feed displaying the environment ahead of the vehicle. In the following subsections we define various terms required to formalize the supervised learning problem.
IV-B1 Trajectory Labels
During the data collection the human operator chooses a control command from $A$ at every instant. The index of this control command uniquely identifies the trajectory/control preference and provides us the ground truth label for that instant. The index label are then converted to one-hot-encoded vectors for the supervised learning training. For an instant $t$, we use $L_{t}$ to represent the corresponding ground truth one-hot-encoded vector.
IV-B2 Prediction
Given the utility feature $U_{t}(S,P)$ at each instant $t$, the classifier outputs the probabilities $F_{t}$ of all the trajectories in the reference set. We use Eq. 2 to calculate the first 4 terrain utility functions. To calculate the last utility function we first transform all the waypoints to the local vehicle frame and then use Eq. 4. We finally use Eq. 1 to calculate the utility feature $U_{t}(S,P)$.
@@ IV-B2 Prediction
Figure 1: Architecture of the classifier. Layers and data are represented by blue and green rectangles respectively. Here $m=21$, Since we have 21 different trajectories in the preference set $S$ (Section V)
@@ IV-B4 Learning objective
V Training and Evaluation
V-A Simulation Setup
Figure 2: shows the AirSim simulation environment with Warthog. The trail shown in this figure is the reference path for training.
We evaluated our algorithm using the AirSim[14] simulation environment. We developed a Warthog unmanned ground vehicle (UGV) model and created an off-road environment with four distinct terrain types: mud, rock, water, and non-traversable. We attach a front facing camera and a GPS sensor on the warthog to get the segmentation images and the vehicle pose respectively. The camera is situated at a height of 1m from the ground and is tilted by 30 degrees downward to get a better view of the terrain in front. Here are some of the details of the evaluation setup:
@@ V-A Simulation Setup
To evaluate our algorithm, we structured the training and testing phases to progressively assess its generalization capabilities. During training, we provided scenarios to the human driver that required making choices between terrain types that are consecutive in Eq. 8. This means the driver only encountered decisions between adjacent terrain preferences in our established hierarchy. In contrast, during testing, the algorithm was presented with scenarios involving terrains without any such restrictions, allowing for choices between non-consecutive terrain types.
@@
Figure 3: Main figure caption for the 1×2 grid of subfigures, spanning both columns.
@@
V-B Training
Fig. 3(a) presents a top-down view of the training environment in AirSim. The reference path for the human operator, shown in white, corresponds to a trail within a grassy landscape that the operator is instructed to follow. This trail includes four terrain patches—water, rock, mud, and non-traversable areas—designed to present different scenarios to the operator. Although the grassy terrain is traversable, in this work we assume that we want to follow the waypoints as closely as possible so the warthog cannot simply take very wide turns to avoid all the terrain types. Note that during the training (Fig. 3(a)), we only provide choices between terrain types that are consecutive in Eq. 8.
Fig. 3(b) shows the actual path navigated by the human operator during training. The scenarios where the operator deviates from the reference path are numbered from 1 to 10. The first four deviations occur in areas containing only a single terrain (no choice), which the operator simply avoids. In contrast, the subsequent six deviations involve scenarios with two or more terrain types, requiring the operator to make choices between them. The various terrains involved in scenarios numbered 1 through 10, along with the corresponding human preferences, are presented in Table I
We collected the training data as described in Section IV-B3 at 10Hz. During the data collection we drive the warthog for approximately 5 mins and collected 2726 samples. This data is then divided into (80,20) split of training examples and validation examples. We then performed 15 epochs of training using the Adam optimizer [15] with a learning rate of $1\times 10^{-3}$ and batch size of 256. Fig. 6 shows the training and validation loss with number of epochs. We save the model checkpoints at each epoch and use the model checkpoint corresponding to the 12th epoch for the testing.
TABLE I: Training scenarios
@@ V-B Training
V-C Testing
Fig. 4 shows the reference path and scenarios for testing. We evaluate our planner on 7 different scenarios that cover a wide variety of terrain configurations in the off-road environment. The first two scenarios have rock and water terrain types with different shape of terrain patches than the ones used in training. The planner trajectory for this scenario is shown in Fig. 5, as we can see from the trajectory that the warthog is able to avoid terrain features of different sizes. The terrain shapes are also kept different from the ones in training data of all the subsequent scenarios. Fig. 5 visualizes the trajectory taken by warthog for all the scenarios and Table-II summarizes the testing results. Fig. 7 and Fig. 8, provide more results with different terrain sizes and scenarios. Fig. 7(a) and Fig. 7(b) show results on terrain with variation in length and width. Fig. 7(c) and Fig. 7(d) shows that the planner can adapt and either go left or right based on where the rocky terrain is. Fig. 8(a) and Fig. 8(b) repeat the same experiment but with different terrain sizes. Fig. 8(c) and Fig. 8(d) show that when presented with 3 different terrains warthog is able to choose the most preferred one (mud) and adapt (going left vs going right). Even with training based on only five minutes of data, these results demonstrate that the planner can adapt to terrain configurations absent from the training set and accommodate variations in terrain size
@@ V-C Testing
TABLE II: Testing Scenarios
@@ VI CONCLUSIONS
The success of our local planner in accurately emulating human driving preferences in off-road scenarios, coupled with its quick adaptability, makes it an attractive candidate for autonomous vehicle navigation in challenging terrains. This research paves the way for more intuitive and human-like autonomous driving solutions but also significantly reduces the barrier to entry for deploying such technologies in diverse off-road environments. Future work will focus on validating the planner’s performance in real-world off-road settings to assess its sim-to-real transfer capabilities along with expanding the range of terrain types, incorporating additional sensor modalities, and integrating with higher-level path planning algorithms to enhance its applicability in real-world autonomous systems.

### v2

@@ Abstract
keywords:
Offroad Robot Navigation, Task and Motion Planning, Path Following
1 INTRODUCTION
@@
Our approach uses utility Dai (2022) functions to directly extract key features from segmented images and learns human driving behaviour using demonstration data. This approach diverges from traditional methods, which typically require either extensive labeled data for end-to-end learning or precise sensor calibration and global mapping in classical robotics approaches. By focusing on extracting key features directly in the trajectory space, our method simplifies the process, avoiding the complexity of global map generation. This allows our algorithm to effectively learn driving patterns from as little as 5-10 minutes of driving data. Our approach reduces the need for large datasets and detailed calibration, facilitating a more straightforward and efficient learning algorithm for navigating complex environments. The main contributions of this work are outlined below:
@@
2 Related Work
Classical work in off-road navigation Kelly (2006); Kim (2007) has focused on creating the costmap of the environment from sensors’ data to represent navigation cost associated with the various types of visual and geometric features of the environment. Earlier approaches Kelly (2006); Thrun (2006) relied on feature engineering while later approaches Maturana (2018); Valada (2017) relied on deep learning based semantic segmentation to represent the visual and geometric features of the surrounding terrain. These later approaches train the semantic segmentation pipeline from scratch and use hand designed cost functions for the planner. Although, these costmaps provide rich information for the downstream planning tasks, tuning them to capture the complex dynamic interactions while navigating on various terrain types is extremely challenging and requires significant domain expertise.
Recent advances in Deep learning have inspired researchers in robotics community to develop end-to-end learning algorithmsBajracharya (2009); Kahn (2020) that directly learn the mapping from sensor information to control commands thereby bypassing the need for manual costmap creation and tuning. Despite the promises of these end-to-end approaches, they still require a large amount of data and show poor generalisation to different settings (domain adaptation). Moreover, the black box nature of these approaches make them very hard to debug and deploy on real systems.
@@ 2 Related Work
Recent works like Triest (2023); Cai (2022) leverage human driving data to either directly learn the costmap of the environment or learn the traversability of various terrain types. These costmaps are crucial to solve the optimal control problem and generate appropriate vehicle controls. While these methods effectively model the tire-terrain interactions of the vehicle, they often overlook the higher level reasoning required to navigate in challenging off-road environments. Additionally, these methods primarily depend on real-world data to learn costmaps and traversability. In contrast, our research demonstrates the potential of simulation platforms to learn a local planning algorithm.
3 Problem Statement
@@ 3.1 Definitions
  • Preference set $S$: The preference set $S$ is a fixed set that represents all the different preferences (Trajectories) available for the human operator at any instant. To construct $S$ we first discretize the control commands—linear and angular velocities. Let $A$, represent this fixed discrete set of linear and angular velocities:
    $A=[\{v^{i},w^{i}\}_{i=1}^{i=m}]$ (1)
    Assuming the vehicle starts at $[0,0,0]$, we then apply forward simulation using the differential drive kinematic model to generate $m$ distinct trajectories over a pre defined time horizon $n$.
@@
    $\theta^{i}_{t}$ $=\theta^{i}_{t-1}+w^{i}\Delta t$
@@ 4 METHODOLOGY
$U(S,P)$ $=\begin{bmatrix}u^{1}(T^{1})&u^{2}(T^{1})&\ldots&u^{5}(T^{1},S,P)\\ u^{1}(T^{2})&u^{2}(T^{3})&\ldots&u^{5}(T^{2},S,P)\\ \vdots&\vdots&\vdots&\vdots\\ u^{1}(T^{m})&u^{2}(T^{m})&\ldots&u^{5}(T^{m},S,P)\end{bmatrix}$ (2)
@@
Here, $u^{1}(T),..,u^{4}(T)$ are utility functions associated with four different terrain features of the environment and depend only on a single trajectory. The utility function $u^{5}(T,S,P)$ represents the distance-based utility, which depends on a trajectory $T$, the preference set $S$ and the reference path $P$. Each row $i$ in Eq. 2 corresponds to all utility functions for the trajectory $T^{i}\in S$. In the next section, we discuss these utility functions in detail.
4.1 Utility Functions
In this work, we consider five utility functions, which are categorized into terrain utility functions and distance utility functions. The first four, $u^{1}(T)$ through $u^{4}(T)$, are terrain utility functions that correspond to different terrain types in the environment: non-traversable, water, rock, and mud, respectively. The fifth utility function, $u^{5}(T,S,P)$,is a distance-based utility function that measures how close a trajectory $T$ is to a reference path $P$, relative to the other trajectories in the preference set $S$. To calculate the utility functions corresponding to the 4 terrain types, we project the trajectory T on to the camera plane and then use the following equation.
@@ 4.1 Utility Functions
The distance utility function is defined with respect to a preference set and a reference path. Given a preference set $S=\{T^{i}\}_{i=1}^{i=m}$, for each trajectory $T^{i}\in S$, let $d(T^{i},P)$ denote the distance between the final point of the trajectory and the nearest waypoint on the reference path $P$. Using this, we define the distance utility function $(u^{5}(T^{i},S,P))$:
@@
Thus the distance utility function represents the relative closeness of a trajectory from the path $P$ within the preference set $S$.
4.2 Learning human preferences using demonstrations
To learn human preferences we frame the problem as a muticriteria decision-making problem where, at each time step, the planner evaluates a set of utilities representing the degree of satisfying multiple criteria (specific terrain preference and closeness to the reference path) and selects a trajectory from a fixed preference set S of trajectories. Since we are dealing with multiple utilities associated with a specific trajectory, we are interested in an overall decision function that aggregates the utilities $(u^{1}(T,S,P),..,u^{5}(T,S,P))$ and indicates the planner’s degree of preference for that possible trajectory. To avoid the cumbersome process of defining the structure and parameters of the decision function Dai (2022) we transform the muticriteria decision-making problem to a classification problem. The labels for this classification task are derived from human demonstrations, where at each time step, the human operator selects a trajectory from the same fixed preference set $S$ while following a reference path. During this demonstration, the human operator follows the reference path as closely as possible, while simultaneously avoiding or preferring different terrain types based on the real-time camera feed displaying the environment ahead of the vehicle. In the following subsections we define various terms required to formalize the supervised learning problem.
4.2.1 Trajectory Labels
During the data collection the human operator chooses a control command from the set $A$ (Eq. 1) at every instant. The index of this control command uniquely identifies the trajectory/control preference and provides us with the ground-truth label for that instant. The index labels are then converted to one-hot-encoded vectors for supervised learning training. For an instant $t$, we use $L_{t}$ to represent the corresponding ground truth one-hot-encoded vector.
4.2.2 Prediction
Given the utility feature $U_{t}(S,P)$ at each instant $t$, the classifier outputs the probabilities $F_{t}$ of all the trajectories in the reference set. We use Eq. 3 to calculate the first 4 terrain utility functions. To calculate the last utility function we first transform all the waypoints to the local vehicle frame and then use Eq. 5. We finally use Eq. 2 to calculate the utility feature $U_{t}(S,P)$. Please note that we only need to compute the preference set $S$ once since the utility feature is calculated in the local vehicle frame at each time instant $t$.
@@ 4.2.2 Prediction
Figure 1: Architecture of the classifier. Layers and data are represented by blue and green rectangles respectively. Here $m=21$, Since we have 21 different trajectories in the preference set $S$ (Section 5)
@@ 4.2.4 Learning objective
5 Training and Evaluation
5.1 Simulation Setup
We evaluated our algorithm using the AirSim Shah (2017) simulation environment. We developed a Warthog unmanned ground vehicle (UGV) model and created an off-road environment with four distinct terrain types: mud, rock, water, and non-traversable. We attach a front facing camera and a GPS sensor on the warthog to get the segmentation images and the vehicle pose respectively. The camera is situated at a height of 1m from the ground and is tilted by 30 degrees downward to get a better view of the terrain in front. Here are some of the details of the evaluation setup:
@@ 5.1 Simulation Setup
To evaluate our algorithm, we structured the training and testing phases to progressively assess its generalization capabilities. During training, we provided scenarios to the human driver that required making choices between terrain types that are consecutive in relation 9. This means the driver only encountered decisions between adjacent terrain preferences in our established hierarchy. In contrast, during testing, the algorithm was presented with scenarios involving terrains without any such restrictions, allowing for choices between non-consecutive terrain types.
@@
Figure 2: Top down view of the AirSim environment
@@
5.2 Training
Fig. 2(a) presents a top-down view of the training environment in AirSim. The reference path for the human operator, shown in white, corresponds to a trail within a grassy landscape that the operator is instructed to follow. This trail includes four terrain patches—water, rock, mud, and non-traversable areas—designed to present different scenarios to the operator. Although the grassy terrain is traversable, in this work we assume that we want to follow the waypoints as closely as possible so the warthog cannot simply take very wide turns to avoid all the terrain types. Note that during the training (Fig. 2(a)), we only provide choices between terrain types that are consecutive in relation 9.
Fig. 2(b) shows the actual path navigated by the human operator during training. The scenarios where the operator deviates from the reference path are numbered from 1 to 10. The first four deviations occur in areas containing only a single terrain (no choice), which the operator simply avoids. In contrast, the subsequent six deviations involve scenarios with two or more terrain types, requiring the operator to make choices between them. The various terrains involved in scenarios numbered 1 through 10, along with the corresponding human preferences, are presented in Table 1
We collected the training data as described in Section 4.2.3 at 10Hz. During the data collection we drive the warthog for approximately 5 mins and collected 2726 samples. This data is then divided into (80,20) split of training examples and validation examples. We then performed 15 epochs of training with a learning rate of $1\times 10^{-3}$ and batch size of 256. Fig. 5 shows the training and validation loss with number of epochs. We save the model checkpoints at each epoch and use the model checkpoint corresponding to the 12th epoch for the testing.
Table 1: Training scenarios
@@ 5.2 Training
5.3 Testing
Fig. 3 shows the reference path and scenarios for testing. We evaluate our planner on 7 different scenarios that cover a wide variety of terrain configurations in the off-road environment. The first two scenarios have rock and water terrain types with different shape of terrain patches than the ones used in training. The planner trajectory for this scenario is shown in Fig. 4, as we can see from the trajectory that the warthog is able to avoid terrain features of different sizes. The terrain shapes are also kept different from the ones in training data of all the subsequent scenarios. Fig. 4 visualizes the trajectory taken by warthog for all the scenarios and Table-2 summarizes the testing results. Fig. 6 and Fig. 7, provide more results with different terrain sizes and scenarios. Fig. 6(a) and Fig. 6(b) show results on terrain with variation in length and width. Fig. 6(c) and Fig. 6(d) shows that the planner can adapt and either go left or right based on where the rocky terrain is. Fig. 7(a) and Fig. 7(b) repeat the same experiment but with different terrain sizes. Fig. 7(c) and Fig. 7(d) show that when presented with 3 different terrains warthog is able to choose the most preferred one (mud) and adapt (going left vs going right). Even with training based on only five minutes of data, these results demonstrate that the planner can adapt to terrain configurations absent from the training set and accommodate variations in terrain size
@@ 5.3 Testing
Table 2: Testing Scenarios
@@ 6 CONCLUSIONS
The success of our local planner in accurately emulating human driving preferences in off-road scenarios, coupled with its quick adaptability, makes it an attractive candidate for autonomous vehicle navigation in challenging terrains. This research paves the way for more intuitive and human-like autonomous driving solutions but also significantly reduces the barrier to entry for deploying such technologies in diverse off-road environments. Future work will focus on validating the planner’s performance in real-world off-road settings to assess its sim-to-real transfer capabilities along with expanding the range of terrain types, incorporating additional sensor modalities, and integrating with higher-level path planning algorithms to enhance its applicability in real-world autonomous systems